# An Investigation of Robustness of LLMs in Mathematical Reasoning: Benchmarking with Mathematically-Equivalent Transformation of Advanced Mathematical Problems

### v2

@@ Abstract
In this paper, we introduce a systematic framework beyond conventional method to assess LLMs’ mathematical-reasoning robustness by stress-testing them on advanced math problems that are mathematically equivalent but with linguistic and parametric variation. These transformations allow us to measure the sensitivity of LLMs to non-mathematical perturbations, thereby enabling a more accurate evaluation of their mathematical reasoning capabilities. Using this new evaluation methodology, we created PutnamGAP, a new benchmark dataset with multiple mathematically-equivalent variations of competition-level math problems. With the new dataset, we evaluate multiple families of representative LLMs and examine their robustness. Across 18 commercial and open-source models we observe sharp performance degradation on the variants. OpenAI’s flagship reasoning model, O3, scores 51.5 % on the originals but drops by 4.7 percentage points on surface-renaming variants, and by 12.9 percentage points on parametric variants, while smaller models fare far worse. Overall, the results show that the proposed new evaluation methodology is effective for deepening our understanding of the robustness of LLMs and generating new insights for further improving their mathematical reasoning capabilities.
@@ 1  Introduction
Benchmark inflation through training leakage. Recent studies show that public datasets, including GSM8K (Cobbe et al., 2021) and MATH (Hendrycks et al., 2021), have leaked into the web‑scale corpora used to pre‑train large language models (LLMs), artificially inflating test‑time accuracy. A leaderboard score therefore no longer guarantees genuine reasoning ability; it may merely reflect memorization of benchmark items or their solutions. Simply releasing yet another dataset postpones the problem: once its items enter future training corpora, scores climb without real progress. What is needed is a systematic method that (i) measures a model’s capacity to generalize beyond verbatim memory and (ii) can generate an unbounded supply of evaluation items, limiting future leakage.
@@
Competition mathematics reveals the next robustness bottleneck. Large language models (LLMs) now surpass 90% accuracy on widely‑used benchmarks such as GSM8K and MATH, prompting claims of “near‑human” numerical reasoning yet still falter on Olympiad‑style or Putnam‑level problems that intertwine multiple domains. Existing Putnam‑derived datasets are too small to expose this gap: Putnam‑AXIOM (236 originals + 52 variations) (Huang et al., 2025), and PutnamBench (640 formalized theorems) (Tsoukalas et al., 2024) remain in the hundreds, and none delivers systematic generalization and perturbations. These facts expose Weakness (i) insufficient scale and Weakness (ii) lack of controlled, systematic transformations in existing evaluations.
Generalization–and–Perturbation (GAP), a novel evaluation strategy. We address both leakage and robustness with a new idea: stress‑test the model on mathematically equivalent versions of the same problem. For a problem $x$ with solution set $S(x)$ and an LLM $f$, robustness is the expected accuracy when $x$ is transformed by a family $\mathcal{T}$ of equivalence‑preserving operators. We partition $\mathcal{T}$ into $\mathcal{T}_{\text{surf}}$ (surface renames that alter symbol salience) and $\mathcal{T}_{\text{para}}$ (kernel rewrites that preserve the same proof steps while changing the scenario and parameters). This GAP framework (i) creates an infinite stream of unseen test items, mitigating future contamination, and (ii) quantifies how far a model can generalize beyond memorized surface forms. It offers a new general diagnostic evaluation methodology for analyzing and quantifying the robustness of an LLM’s mathematical reasoning capacity.
@@
Contributions. (1) We propose GAP, a novel general framework for measuring robustness via mathematically equivalent transformations that overcomes two common deficiencies of the current evaluation methods (i.e., data leakage and lack of robustness measures). (2) We release PutnamGAP, the first 6k‑scale competition benchmark that systematically disentangles surface‑level and structural generalization while limiting future leakage. (3) We provide the first comprehensive robustness baseline across seventeen LLMs, plus an open‑source evaluation stack.
@@ 5.1 Robustness
We evaluated 18 different LLMs on this benchmark, and results are summarized in Table 1. For each variation of the model, we used a paired design (McNemar’s exact test) on matched problem pairs to test whether the accuracy rate decreases significantly compared to the original. Statistically significant differences are indicated using standard notation ($p<0.1$, $p<0.05$, $p<0.01$). We also computed 95 % CI (See Appendix D Figure 4 ) and robustness metrics $R$ (See Appendix D Figure 9 ), and all models, especially those performed well on the original set, got a low robustness score.
@@
[…]
Another observation is if a model is not robust on one variant, it tends to be not robust on other variants. Notable examples would be kimi-k2, cluase-opus-4 and gemini-2.5-pro.
@@ 5.4 External Validation
We applied our surface-renaming protocols—DLC and GS—to ALG514  (kus, 2014). Accuracy decreased from Base 93.6% to DLC 90.9% ($\Delta=-2.7$ pp) and GS 89.3% ($\Delta=-4.3$ pp); McNemar tests (Base vs DLC: b=24,c=10,p=0.024b{=}24,c{=}10,p{=}0.024; Base vs GS: b=35,c=13,p=0.002b{=}35,c{=}13,p{=}0.002). These statistically significant drops indicate that GAP’s surface-renaming stress tests generalize to other math datasets and reveal nontrivial sensitivity to variable renaming.
@@ 6.1 Key Findings
Maintaining structure but resampling parameters is even harsher. The Kernel Variant (KV) simultaneously resamples all mutable constants while preserving the original reasoning skeleton. Accuracy losses reach $\approx 10$ pp; OpenAI O3 declines from 48.8% to 38.5% (–10.3 pp), showing that grasping a solution pattern does not automatically translate to parameter-invariant reasoning ability.
@@ 7  Related Work
There have been multiple benchmarks for evaluating the mathematical-reasoning capabilities of large language models (LLMs). Early math‐reasoning benchmarks such as MATH(1.25 k problems) (Hendrycks et al., 2021), and GSM8K(8.5 k problems) (Cobbe et al., 2021), revealed basic arithmetic/algebra skills. But their difficulty is now saturated as LLMs scale. For instance, with prompting strategies such as DUP, GPT-4 attains 97.1% accuracy on GSM8K (Zhong et al., 2025). This ceiling at the high-school-competition level motivated the creation of a new generation of harder benchmarks.
@@
Subsequent benchmarks target harder problems. Omni-MATH contributes 4 428 rigorously annotated Olympiad-level problems (Gao et al., 2024). Likewise, OlympiadBench provides a bilingual, multimodal benchmark of 8 476 Olympiad-level math and physics problems with expert step-by-step solutions (He et al., 2024). The cross-disciplinary benchmark ARB consist questions in mathematics, physics, biology, chemistry, and law, with a rubric-based self-grading protocol (Sawada et al., 2023). Some other benchmarks focuses specifically on formal proof. MiniF2F supplies 488 Olympiad-level problems formalized in multiple proof assistants (Zheng et al., 2022). Putnambench, offers 1 692 rigorously hand-crafted formalizations of Putnam Competition problems (Tsoukalas et al., 2024).
@@
Nevertheless, recent studies warn that scores on many NLP benchmarks may be artificially inflated by data contamination, when LLMs are trained on the benchmark questions. Sainz et al. (2023) point out that many benchmarks may be inflated because large language models often memorize test data seen during pre-training.  Balloccu et al. (2024) conduct a systematic audit of data leakage for closed-source LLMs and estimate that roughly 4.7 million test examples from 263 datasets were likely exposed to the models.
In order to obtain a more robust evaluation of LLMs’ reasoning capabilities, it is important to prevent data leakage. One approach is to create original questions. FrontierMath, for example, addresses this with a rigorously curated benchmark of hundreds of original, expert-level mathematics problems spanning fields from number theory to algebraic geometry (Glazer et al., 2024). Putnam-AXIOM takes this approach and comprises 522 challenging problems from the William Lowell Putnam Competition plus 100 programmatically generated functional variations, furnishing a contamination-resilient benchmark (Gulati et al., 2025).
Another method to deal with data leakage is introducing contrast sets-small, label-changing perturbations of existing test instances—to probe a model’s local decision boundary (Gardner et al., 2020). Huang et al. (2025) construct MATH-Perturb that applies simple and hard perturbations to 279 level-5 MATH problems and discover that models suffer 12–16 percentage-point drops on the hard variant. Shalyt et al. (2025) complement this line of work with ASyMOB, a 17k-problem benchmark whose algebra-focused numerical and symbolic perturbations reveal performance drops of up to 70 percentage points, highlighting models’ fragility under such stress tests. Similarly, Yu et al. (2025) propose Math-RoB, a synthetic benchmark that enables robustness evaluation against the data contamination through an instruction-based approach. These efforts either focus on specific aspects that limit generalizability, are based on benchmarks that are too easy for current models, or introduce transformations that are not mathematically equivalent, thus confounding true robustness evaluation.
@@
Building on these prior efforts, our work introduces Generalization–and–Perturbation (GAP), a unified framework that addresses both data leakage and robustness by generating mathematically equivalent variants of complex problems, significantly expanding the evaluation depth of existing benchmarks. The framework can be applied to existing and future benchmarks, and all types of questions to strengthen their reliability. To address the saturation of accuracy scores, we apply our framework to challenging, college-level competition mathematics problems. We instantiate GAP on every William Lowell Putnam Competition problem from 1938–2024 (1 051 originals), expanding each item into five mathematically equivalent variants and thereby producing PutnamGAP, a corpus of 6,306 stress-test questions. Finally, we release an open-source evaluation stack that rigorously grades solutions step by step, making assessment fully automated, transparent, and reproducible.

### v3

@@ Abstract
In this paper, we introduce a systematic framework beyond conventional methods to assess LLMs’ mathematical‑reasoning robustness by stress‑testing them on advanced math problems that are mathematically equivalent but with linguistic and parametric variation. These transformations allow us to measure the sensitivity of LLMs to non-mathematical perturbations, thereby enabling a more accurate evaluation of their mathematical reasoning capabilities. Using this new evaluation methodology, we created PutnamGAP, a new benchmark dataset with multiple mathematically-equivalent variations of competition-level math problems. With the new dataset, we evaluate multiple families of representative LLMs and examine their robustness. Across 18 commercial and open-source models we observe sharp performance degradation on the variants. OpenAI’s flagship reasoning model, O3, scores 51.5 % on the originals but drops by 4.7 percentage points on surface-renaming variants, and by 12.9 percentage points on parametric variants, while smaller models fare far worse. Overall, the results show that the proposed new evaluation methodology is effective for deepening our understanding of the robustness of LLMs and generating new insights for further improving their mathematical reasoning capabilities.
@@ 1  Introduction
Benchmark inflation through training leakage. Recent studies show that public datasets, including GSM8K (Cobbe:2021) and MATH (hendrycks:2021), have leaked into the web‑scale corpora used to pre‑train large language models (LLMs), artificially inflating test‑time accuracy. A leaderboard score therefore no longer guarantees genuine reasoning ability; it may merely reflect memorization of benchmark items or their solutions. Simply releasing yet another dataset postpones the problem: once its items enter future training corpora, scores climb without real progress. What is needed is a systematic method that (i) measures a model’s capacity to generalize beyond verbatim memory and (ii) can generate an unbounded supply of evaluation items, limiting future leakage.
@@
Competition mathematics reveals the next robustness bottleneck. Large language models (LLMs) now surpass 90% accuracy on widely‑used benchmarks such as GSM8K and MATH, prompting claims of “near‑human” numerical reasoning yet still falter on Olympiad‑style or Putnam‑level problems that intertwine multiple domains. Existing Putnam‑derived datasets are too small to expose this gap: Putnam‑AXIOM (236 originals + 52 variations) (huang:2025), and PutnamBench (640 formalized theorems) (tsoukalas:2024) remain in the hundreds, and none delivers systematic generalization and perturbations. These facts expose Weakness (i) insufficient scale and Weakness (ii) lack of controlled, systematic transformations in existing evaluations.
Existing perturbation-based robustness benchmarks. Recent work has begun to probe mathematical robustness by constructing perturbation-based benchmarks on top of GSM8K and related datasets. GSM-Plus augments GSM8K with eight families of adversarial variations per problem, revealing large accuracy drops even for models that nearly solve the original benchmark (Li:2024GSMPlus). GSM-Symbolic builds symbolic templates over GSM8K-style problems and shows that merely changing numeric instantiations or adding logically irrelevant clauses can degrade performance by up to 65% (Mirzadeh:2024GSMSymbolic). MathCheck-GSM further organizes GSM8K-derived problems into a checklist of task and robustness variants to study behavior across multiple evaluation formats (Zhou:2024MathCheck). Beyond GSM8K, GSM8K_MORE uses an ontology of perturbations to generate families of grade-school arithmetic variants (Hong:2025GSMore), while Putnam-AXIOM introduces a smaller set of functional variations for university-level Putnam problems (gulati:2025). These efforts convincingly demonstrate that current LLMs are brittle under controlled perturbations; however, GSM-derived benchmarks remain confined to grade-school or pre-university word problems with short, single-answer numerical solutions and are built directly on GSM8K and related datasets that are already near-saturated and affected by training data contamination for frontier models (Cobbe:2021; gulati:2025; Shalyt:2025; glazer:2024), while Putnam-AXIOM introduces only a relatively small companion set of functional variants (100 over 522 problems) . Consequently, the existing perturbation benchmarks do not yet provide a large-scale, systematically structured robustness test for competition-level, proof-style mathematics.
Generalization–and–Perturbation (GAP) framework for robustness evaluation.
We address both leakage and robustness by stress-testing the model on mathematically equivalent versions of the same problem. For a problem $x$ with solution set $S(x)$ and an LLM $f$, robustness is the expected accuracy when $x$ is transformed by a family $\mathcal{T}$ of equivalence-preserving operators. We partition $\mathcal{T}$ into $\mathcal{T}_{\text{surf}}$ (surface renames that alter symbol salience) and $\mathcal{T}_{\text{para}}$ (kernel rewrites that preserve the same proof steps while changing the scenario and parameters). This GAP framework (i) creates an infinite stream of unseen test items, mitigating future contamination, and (ii) quantifies how far a model can generalize beyond memorized surface forms. In our setting, GAP serves as a general diagnostic evaluation methodology for analyzing and quantifying the robustness of an LLM’s mathematical reasoning capacity at the level of competition problems.
Limitations of existing perturbation benchmarks. Several recent robustness benchmarks - such as GSM-Symbolic, GSM-Plus, and MathCheck -
@@
Contributions. (1) We propose GAP, a novel general framework for measuring robustness via mathematically equivalent transformations that overcomes two common deficiencies of the current evaluation methods (i.e., data leakage and lack of robustness measures). (2) We release PutnamGAP, the first 6k‑scale competition benchmark that systematically disentangles surface‑level and structural generalization while limiting future leakage. (3) We provide the first comprehensive robustness baseline across eighteen LLMs, plus an open‑source evaluation stack.
@@ 5.1 Robustness
We evaluated 18 different LLMs on this benchmark, and results are summarized in Table 1. For each variation of the model, we used a paired design (McNemar’s exact test) on matched problem pairs to test whether the accuracy rate decreases significantly compared to the original. Statistically significant differences are indicated using standard notation ($p<0.1$, $p<0.05$, $p<0.01$). We also computed 95 % CI (See Appendix D Figure 4 ) and our proposed robustness metrics $R$ (see Appendix B), and all models, especially those performed well on the original set.
@@
Because the surface score aggregates the four renaming variants by per–item majority, the flip probability from the original to the aggregated surface set is suppressed; accordingly, $R\!\approx\!1$ is expected and should be interpreted as an approximate upper bound on surface invariance (see Table 2). Practitioners can implement alternative mapping functions based on their model’s performance while retaining this core formulation. Across capable models we consistently observe $R_{\text{para}}<R_{\text{surf}}$, and we summarize stress–type invariance via $R_{\text{global}}=\sqrt{R_{\text{surf}}R_{\text{para}}}$. Interpreting 1R1-R as a penalty mass highlights nontrivial fragility even when raw accuracy is high. Conversely, for weak models a high $R$ is not evidence of robustness: when base accuracy $p_{e}$ is small, the pooled SD $\sigma=\sqrt{\tfrac{1}{2}\!\left(p_{e}(1-p_{e})+p_{h}(1-p_{h})\right)}$ and the bound $1-R\leq\min\{p_{e},1-p_{h}\}\,(1-q)$ with $q=\exp(-\beta d_{b+})$ limit the observable penalty, so $R\to 1$ reflects low headroom rather than invariance. Reporting both accuracy and $\{R_{\text{surf}},R_{\text{para}},R_{\text{global}}\}$ therefore stabilizes cross–model comparison under mathematically equivalent stress and shows that robustness remains limited despite strong performance on canonical phrasing.
Another observation is that if a model is not robust to one variant, it tends to be not robust to other variants as well. Notable examples include kimi-k2, claude-opus-4, and gemini-2.5-pro.
@@ 5.3 Error Taxonomy
5.4 Qualitative case studies of Kernel Variant failures
To complement the aggregate robustness metrics, we performed a small-scale qualitative analysis of Kernel Variant (KV) failures. We ran a GPT-based analyzer over model traces and automatically selected ORIGINAL/KV pairs where a strong model solves the ORIGINAL correctly but fails on the KERNEL-VARIANT; concrete case studies are deferred to Appendix I.
Across these examples we see three recurring KV-specific failure modes. First, hallucinated algebraic infrastructure and missing premises: in items such as 1938-B-1 and 1940-A-6 the KV solutions invoke strong algebraic identities or valuation equalities (e.g., $\operatorname{adj}M=(\det M)M^{-1}$ or $v_{i}(JF)=e_{i}-1$) without checking that the hypotheses hold in the stated ring or characteristic, whereas the ORIGINAL proofs stay within a valid algebraic framework. Second, computing the wrong global quantity after mostly correct setup: in 1939-A-1, 1940-A-7, and 1940-B-7 the KV traces correctly identify the relevant points or bounds but then switch from arc length to chord length or from a clean monotonicity argument to a mis-indexed summation, producing false inequalities despite reasonable intermediate calculus or algebra. Third, fragile geometric reductions and inconsistent conventions: in 1939-B-1, 1939-B-7, 1940-A-2, and 1938-A-7 the KV arguments rely on incorrect symmetry reductions, ignore degenerate edge cases (e.g. $\rho=0$), or briefly adopt sign conventions that contradict earlier definitions before silently reverting.
Overall, these qualitative patterns corroborate the quantitative gap $R_{\text{para}}<R_{\text{surf}}$. Kernel Variants do not merely inject harder arithmetic; they stress the model’s ability to re-bind parameters and maintain a coherent proof skeleton under resampled slots. When the model fails KV, it often does so by reusing an ORIGINAL template outside its domain of validity or by quietly changing the quantity or symmetry being computed (see Appendix I for detailed traces).
@@ 5.5 External Validation
We applied our surface-renaming protocols—DLC and GS—to ALG514  (kushman-etal-2014-learning). Accuracy decreased from Base 93.6% to DLC 90.9% ($\Delta=-2.7$ pp) and GS 89.3% ($\Delta=-4.3$ pp); McNemar tests (Base vs DLC: b=24,c=10,p=0.024b{=}24,c{=}10,p{=}0.024; Base vs GS: b=35,c=13,p=0.002b{=}35,c{=}13,p{=}0.002). These statistically significant drops indicate that GAP’s surface-renaming stress tests generalize to other math datasets and reveal nontrivial sensitivity to variable renaming.
@@ 6.1 Key Findings
Maintaining structure but resampling parameters is even harsher. The Kernel Variant (KV) simultaneously resamples all mutable constants while preserving the original reasoning skeleton. Accuracy losses reach $\approx 10$ pp; OpenAI O3 declines from 51.5% to 38.6% (–12.9 pp), showing that grasping a solution pattern does not automatically translate to parameter-invariant reasoning ability.
@@ 7  Related Work
There have been multiple benchmarks for evaluating the mathematical-reasoning capabilities of large language models (LLMs). Early math‐reasoning benchmarks such as MATH(1.25 k problems) (hendrycks:2021), and GSM8K(8.5 k problems) (Cobbe:2021), revealed basic arithmetic/algebra skills. But their difficulty is now saturated as LLMs scale. For instance, with prompting strategies such as DUP, GPT-4 attains 97.1% accuracy on GSM8K (zhong:2025). This ceiling at the high-school-competition level motivated the creation of a new generation of harder benchmarks.
@@
Subsequent benchmarks target harder problems. Omni-MATH contributes 4 428 rigorously annotated Olympiad-level problems (gao:2024). Likewise, OlympiadBench provides a bilingual, multimodal benchmark of 8 476 Olympiad-level math and physics problems with expert step-by-step solutions (he:2024). The cross-disciplinary benchmark ARB consist questions in mathematics, physics, biology, chemistry, and law, with a rubric-based self-grading protocol (sawada:2023). Some other benchmarks focuses specifically on formal proof. MiniF2F supplies 488 Olympiad-level problems formalized in multiple proof assistants (zheng:2022). Putnambench, offers 1 692 rigorously hand-crafted formalizations of Putnam Competition problems (tsoukalas:2024).
@@
Nevertheless, recent studies warn that scores on many NLP benchmarks may be artificially inflated by data contamination, when LLMs are trained on the benchmark questions. sainz:2023 point out that many benchmarks may be inflated because large language models often memorize test data seen during pre-training.  balloccu:2024 conduct a systematic audit of data leakage for closed-source LLMs and estimate that roughly 4.7 million test examples from 263 datasets were likely exposed to the models.
Preventing data leakage is central to obtaining a robust evaluation of LLMs’ reasoning capabilities. One approach is to construct entirely original problems: for example, FrontierMath provides a rigorously curated benchmark of hundreds of original, expert-level mathematics problems spanning fields from number theory to algebraic geometry (glazer:2024). Another strategy is to introduce contrast sets—small, label-changing perturbations of existing test instances—to probe a model’s local decision boundary (gardner:2020). Within this perturbation paradigm, GSM-Plus, GSM-Symbolic, MathCheck-GSM, and GSM8K_MORE all build on GSM8K (Cobbe:2021), augmenting grade-school word problems with adversarial numeric, lexical, and contextual variations and revealing substantial robustness failures (Li:2024GSMPlus; Mirzadeh:2024GSMSymbolic; Zhou:2024MathCheck; Hong:2025GSMore). At higher difficulty, huang:2025 construct MATH-Perturb by applying simple and hard perturbations to 279 level-5 MATH problems, Shalyt:2025 introduce ASyMOB, a 17k-problem benchmark focused on algebraic symbolic operations with numerical and symbolic perturbations, yu:2025 propose Math-RoB, a synthetic benchmark that uses instruction-based modifications to expose reasoning gaps under data contamination, and Putnam-AXIOM combines 522 original Putnam problems with 100 functional variants obtained by perturbing variables and constants (gulati:2025). Collectively, these benchmarks demonstrate that current LLMs are far from robust, but GSM-based variants remain at grade-school arithmetic level on benchmarks that are increasingly saturated and contaminated for frontier models (Cobbe:2021; gulati:2025; Shalyt:2025; glazer:2024), MATH-Perturb and ASyMOB target relatively narrow slices of mathematics (hard MATH items and symbolic algebra, respectively), Math-RoB relies on synthetic instruction-style perturbations that are not strictly mathematically equivalent, and existing Putnam variants form only a small companion set to the original (potentially contaminated) problems.
@@
Building on these prior efforts, we adopt a Generalization–and–Perturbation (GAP) framework that addresses both data leakage and robustness by generating mathematically equivalent variants of complex problems and jointly evaluating models on originals and variants. The framework is agnostic to any particular dataset and can in principle be applied to existing and future benchmarks, and to both proof-style and short-answer questions, to strengthen their reliability. To move beyond saturated, pre-university settings, we apply GAP to challenging college-level competition mathematics problems. Concretely, we instantiate GAP on every William Lowell Putnam Competition problem from 1938–2024 (1 051 originals), expanding each item into five mathematically equivalent variants and thereby producing PutnamGAP, a corpus of 6 306 stress-test questions. Finally, we release an open-source evaluation stack that rigorously grades solutions step by step, making assessment fully automated, transparent, and reproducible.